OpenStudy (anonymous):

find the rank of the matrix 3 0 2 2 -6 42 24 54 21 -21 0 -15

OpenStudy (anonymous):

i.e., the vector potential is: \[ \vec{F}=(y+z)\hat{i}+(z+x)\hat{j}+(x+y)\hat{k} \] fot the line integral to be path independent, we show that F is conservative. i.e.,\[\nabla\times\vec{F}=0\\ \left|\begin{matrix}\hat{i}&\hat{j}&\hat{k}\\ {\partial\over\partial x}&{\partial\over\partial y}&{\partial\over\partial z}\\ y+z&z+x&x+y \end{matrix}\right|=[1-1]\hat{i}-[1-1]\hat{j}+[1-1]\hat{k}=\vec{0} \] this, F is conservative and independent of path

OpenStudy (anonymous):

thanks for solution.....